OpenStudy (anonymous):

You have an SRS of size n = 9 from a normal distribution with sigma = 1 . You wish to test H0: u=0 and Ha: u > 0 You decide to reject Ho 0 if x > 0 and to accept Ho otherwise. a) Find the probability of a Type I error. b) Describe a Type II error.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

what is H0?

OpenStudy (anonymous):

Ho is the null

OpenStudy (michele_laino):

I'm sorry, I don't know your answer

OpenStudy (anonymous):

do you know who to calculate probability of type I error?

OpenStudy (michele_laino):

is it the gaussian probability?

OpenStudy (anonymous):

i dont know

OpenStudy (michele_laino):

n is the degree of freedom?

OpenStudy (anonymous):

umm i don't know, this question is under types of error and powers, if that helps at all

OpenStudy (michele_laino):

please, wait I ask to another tutor

OpenStudy (michele_laino):

@dan815 can you help here please?

OpenStudy (michele_laino):

@sleepyjess please can you help her?

OpenStudy (michele_laino):

here*

OpenStudy (michele_laino):

@freckles can you help here, please?

OpenStudy (anonymous):

thank you @Michele_Laino

OpenStudy (michele_laino):

thank you! Nevertheless your question is without answer

OpenStudy (anonymous):

idk what to do :(

OpenStudy (michele_laino):

what subject are you studying at the moment?

OpenStudy (anonymous):

ap statistics

OpenStudy (michele_laino):

what is the meaning of SRS?

OpenStudy (anonymous):

simple random sample

OpenStudy (michele_laino):

please wait, I'm looking for your answer in my textbook

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

I found this formula: \[\Large \sigma = \sqrt {\frac{N}{{12}}} \] where N is the size of your sample, and sigma is the requested error

OpenStudy (anonymous):

so in our case n is 9 so we would do sqrt(9/12)?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

ok i get 0.86602540378

OpenStudy (michele_laino):

ok! better is 0.9

OpenStudy (anonymous):

ok so thats the answer for a?

OpenStudy (michele_laino):

yes, I think so!

OpenStudy (michele_laino):

no, please wait

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

I think that the requested probability, is given by the subsequent formula: \[\Large p = \frac{{\sqrt {\frac{N}{{12}}} }}{{{\sigma _0}}}\] where: \[{{\sigma _0}}\] is your sigma,namely: \[{{\sigma _0} = 1}\]

OpenStudy (anonymous):

so (sqrt(9/12))/sigmao = .9/1 = .9

OpenStudy (michele_laino):

yes! sorry better is 87%

OpenStudy (michele_laino):

since we got 0.866....

OpenStudy (anonymous):

oh okay so for a we would get .87

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

ok and for b do i just say a false null hypothesis can fail to be rejected

OpenStudy (michele_laino):

wait a moment, I'm reading my textbook

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

yes! I think so, since we got this interval: \[\Large \left( { - 0.87 \times \sqrt {27} ,\;0.87 \times \sqrt {27} } \right) = \left( { - 4.52{\sigma _0},\;4.52{\sigma _0}} \right)\]

OpenStudy (michele_laino):

the general formula which gives the amplitude of that interval is: \[\Large {\frac{N}{{2}}{\sigma _0}}\] @dennese13

OpenStudy (anonymous):

so 3.5?

OpenStudy (anonymous):

*4.5

OpenStudy (michele_laino):

yes! \[\Large {4.52{\sigma _0}}\] is also a right answer

OpenStudy (anonymous):

wait so what would that go under

OpenStudy (michele_laino):

are you referring to the underlying theory?

OpenStudy (anonymous):

no like would that go under b?

OpenStudy (michele_laino):

I think what you have said before, namely: "a false null hypothesis can fail to be rejected"

OpenStudy (anonymous):

wait so where do we putt he 4.5

OpenStudy (michele_laino):

I have provided that number, in order to show you that the probability of part A, is very high. Namely, we need of an objective criterion in order to answer to part B

OpenStudy (michele_laino):

The formulas which I provided you are from the theory. If you want I give you the title of my textbook

OpenStudy (anonymous):

sure!

OpenStudy (anonymous):

so for my answer I can put a) .87 b) false null hypothesis can fail to be rejected

OpenStudy (michele_laino):

yes! here is the title: \[\large \begin{gathered} {\mathbf{Data\;reduction\; and\; error \; analysis\; for \; the\; physical \; sciences}} \hfill \\\; {\text{Authors: P}}{\text{.R}}{\text{.Bevington}}{\text{, D}}{\text{.K}}{\text{. Robinson}} \hfill \\ {\text{Publisher: McGraw - Hill}} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

okay thank you so much!!

OpenStudy (michele_laino):

Thank you!!

OpenStudy (michele_laino):

oops.. the last word of the title, is: \[\Large {\mathbf{sciences}}\]